1) A tank 16m long and 23m wide contains water. how many cubic meters of water must be rushed into it to make the surface rise by 16 2/3 cm?
2) the areas of 3 adjacent faces of a rectangular box which meet in a point are known. The product of these areas is equal to ....................
3) A room 3.3.m high is half as long as it is wide and its volume is 123 3/4 cubic meters. find its length and breadth.

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

length = 16 m
width = 23 m
height = 16 2/3 cm = 16 + 0.66 cm ------ = 16.666 cm -> 0.166 m
Volume of water required to raise the level = 16 * 23* 0.166
= 16.088 m^3

Answer by ikleyn(53618)

.
1) A tank 16m long and 23m wide contains water. how many cubic meters of water must be rushed into it
to make the surface rise by 16 2/3 cm?
~~~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth for problem (1) are incorrect.
        I came to bring a correct solution.


length = 16 m
width = 23 m
height = 16 2/3 cm is about 16 + 0.666 cm, or 16.666 cm, or 0.1666 m
Volume of water required to raise the level = 16 * 23* 0.1666
= 61.088 m^3